OpenStudy (anonymous):

The value of Goodstudent stock increased from $300 a share to $600 a share. What was the percent of increase? 30% 60% 90% 100% 200%

OpenStudy (anonymous):

600 - 300 = 300, 300/300 = 1 *100% = 100%